OpenStudy (anonymous):

There are two candles with different heights and thicknesses.The shorter one lasts 11 hours and the longer one lasts 7 hours when both are lit simontaneously . If after 3 hours of their being lit together at the same time , both have same length , then ratio of their original lengths is ?

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

spouse that X is the length of the tall one. and Y is the length of the short one. then : (4*x)/7 = (8*y)/11 => y=0.78*x

OpenStudy (anonymous):

suppose candle to be cylindrical the options are 5:8 11:14 10:13 7:11

OpenStudy (anonymous):

there is thickness also

OpenStudy (anonymous):

and this ques is supposed to be for 8th class LOL!

OpenStudy (anonymous):

ok see : I said that y=0.7857142857*x then if x=14 (the second option) => y= 0.7857142857 * 14=11

OpenStudy (anonymous):

this is the right answer : 11:14 you dont have to use thickness to solve this problem

OpenStudy (anonymous):

WOW